# Deep Learning Framework Testing via Heuristic Guidance Based on Multiple Model Measurements

### v1

@@ 5.3 RQ4: Effectiveness of Bug Detection
5.3.1 Crashes
As introduced in Section 4.5, DLMMM records the logs while executing test input models to detect crashes. Since one crash may be repeatedly triggered on different datasets, two researchers analyze the logs respectively to remove redundant crashes. Table IV shows the crashes detected by DLMMM under each dataset. In this table, the first column represents the adopted dataset, and the following three columns represent the number of crashes detected under TensorFlow, PyTorch, and MindSpore. Figure 2 shows the total number of crashes triggered on all datasets after removing redundancy. In this figure, the first column represents the method name, and the following three columns represent the number of crashes detected under each tested DL framework. The experimental results show that DLMMM newly detects 19 crashes, including 3 in TensorFlow, 3 in PyTorch and 13 in MindSpore, which is the most among the four baselines. Specifically, only one crash is triggered by LEMON; 13 are by COMET, and 11 are by Gandalf. Muffin does not trigger any crashes. We have reported all the detected crashes to the public community. Among them, 16 of 19 have been confirmed by developers and two of 19 are fixed in the updated version.
In addition, we analyze the root causes behind detected crashes. In TensorFlow, one crash arises from failure in initialization. Another two crashes are due to unsupported implementation (e.g., unsupported operator $sparse\_mat\_mul$ and unsupported parameter $dilations$ for the operator $Convolution$). In PyTorch, one crash arises from the failure in creating a descriptor. Another two crashes are due to unsupported implementation. In MindSpore, one crash arises from unsupported implementation in the operator $PRELU$ in the CPU environment. One crash arises from the failure in creating a descriptor. Another 11 crashes are caused by the failure in memory allocation. This issue leads to crashes during executing multiple operators, such as $Convolution$ and $Depthwise\_Convolution$.
TABLE IV: Number of Unique Crashes Detected by DLMMM
@@ 5.3.1 Crashes
Figure 2: Number of Unique Crashes Detected by Each Method
5.3.2 NaN & Inconsistency Bugs
Table V shows the number of NaN & inconsistency bugs triggered on all datasets. In this table, the first column represents the adopted dataset, and the following three columns represent the number of NaN & Inconsistency bugs detected under each tested DL framework. This table shows NaN & inconsistency bug is successfully triggered in almost all datasets (except MNIST). The small test input tensor size in MNIST (only 48 in height and width dimension) limits the bug detection effectiveness. Figure 3 shows the number of unique NaN & inconsistency bugs detected by DLMMM and all baselines. In this figure, the first column represents the method name, the following three columns represent the same meaning as Table V. The experimental results show that DLMMM triggers most NaN & inconsistency bugs compared to four baselines. Specifically, DLMMM triggers 175 NaN & inconsistency bugs in total, including 73 under TensorFlow, 60 under PyTorch, and 42 under MindSpore. LEMON and Muffin do not detect any NaN & inconsistency bugs at all. COMET and Gandalf detect six and 36 NaN & inconsistency bugs, respectively, under all these frameworks. In conclusion, DLMMM performs best in detecting NaN & inconsistency bugs compared to all baselines under all frameworks.
@@ 5.3.2 NaN & Inconsistency Bugs
TABLE V: Number of Unique NaN & Inconsistency Bugs Detected by DLMMM
@@
Figure 3: Number of Unique Detected NaN & Inconsistency Bugs
@@
5.4 RQ5: Effectiveness of Test Input Model Generation
@@
5.4.1 Operator Combination Variety
Figure 4 shows the average operator combination variety of DLMMM on each dataset. In this figure, the text on the axis represents the datasets, and the numbers on the bars represent the average operator combination variety of each dataset (measured by $variety\_degree$ introduced in Section 3.2.2). Figure 5 shows the comparison between DLMMM and baselines in average operator combination variety in all datasets. In this figure, the text on the axis represents the method name, and the numbers on the bars represent the average operator combination variety of each method. The experimental results show that the average operator combination variety of test input models generated by DLMMM is 64.25, which greatly exceeds 7.98 in LEMON, 5.51 in Muffin, 18.35 in COMET, and 2.56 in Gandalf. Compared to the best performance of all baselines, DLMMM achieves a 250.13% improvement in operator combination variety.
Figure 4: Operator Combination Variety of DLMMM
Figure 5: Comparison between DLMMM and Baselines in Operator Combination Variety
5.4.2 Model Execution Time (Efficiency)
To evaluate the efficiency of DLMMM, we count the number of models executed by DLMMM in six hours and calculate the average model execution time. The experimental results are shown in Table VI. In this table, the first column represents the adopted dataset. The second column represents the average model execution time. The last column represents the number of executed test input models. In addition, we calculate the average number of executed test input models on all datasets and compare it with four baselines, as shown in Figure 6. In this figure, the first column represents the method name, and the meaning of the following two columns is the same as in Table VI. The experimental results show that, under all datasets, DLMMM takes 5.06 seconds at most and 3.50 seconds on average to execute each single model. However, LEMON, Muffin, COMET, and Gandalf take 63.90 seconds, 14.46 seconds, 122.72 seconds, and 36.92 seconds, respectively. In conclusion, the time consumption of DLMMM to execute a single model is within 35% of all state-of-the-art methods.
TABLE VI: Efficiency of DLMMM
@@ 5.4.2 Model Execution Time (Efficiency)
Figure 6: Comparison between DLMMM and Baselines in Efficiency
@@
5.5 RQ6: Ablation Study on Heuristic Guidance
@@
We respectively run DLMMM and all baselines for six hours and record detected crashes and NaN & inconsistency bugs. For fairness, all test input tensors are randomly generated and kept the same among DLMMM and all baselines. The experimental results are shown in Table VII. In this table, the first column represents the method name. The following two columns represent the number of detected crashes and NaN & inconsistency bugs. From the experimental results, we have the following conclusions. Firstly, DLMMM (detecting 19 crashes and 175 NaN & inconsistency bugs) detects more bugs than $DLMMM_{r}$ (detecting no crash and one NaN & inconsistency bug), which shows the positive contribution of our heuristic guidance. Secondly, DLMMM detects more bugs than $DLMMM_{p}$ (detecting one crash and eight NaN & inconsistency bugs), $DLMMM_{o}$ (detecting five crashes and two NaN & inconsistency bugs), and $DLMMM_{t}$ (detecting one crash and three NaN & inconsistency bugs), which shows the positive contribution of our measurement fusion. Thirdly, $DLMMM_{p}$, $DLMMM_{o}$ and $DLMMM_{t}$ all detect more bugs than $DLMMM_{r}$, which respectively shows the positive contribution of bug detection effectiveness, operator combination variety, and model execution time.
TABLE VII: Comparison in Effectiveness in Ablation Study
@@
Answer to RQ6: The ablation study shows that our heuristic guidance, operator fusion, and each model measurement all contribute to improving DLMMM’s bug detection effectiveness.
6 Validity & Threat
@@
The external threat lies in the DL frameworks tested. To eliminate this threat, we apply our method to test three widely used frameworks, respectively TensorFlow, PyTorch and MindSpore. The other frameworks can also be tested by applying the core idea of our method.
@@ 7 Related Works
DL framework bugs have received extensive attention in recent years. Many empirical studies on DL framework bugs are conducted, which establishes the foundation for DL framework testing [25, 26, 27, 28, 29, 30, 31, 32, 33]. DL framework testing basically contains two types: interface-based DL framework testing and model-based DL framework testing. Interface-based DL framework testing separately invokes each operator provided by DL frameworks. Among them, Predoo [20] mutates test input tensors to maximize output precision error. FreeFuzz [34] conducts automated API-level fuzzing via mining from actual code and model executions in the wild. DocTer [35] analyzes API documentation to extract input constraints for each API function provided by DL frameworks. DeepREL [36] infers potential API relations based on their syntactic or semantic information, and synthesizes test programs for invoking relational APIs. Duo [37] generates test input tensors by employing nine mutation rules derived from genetic algorithms. EAGLE [38] invokes equivalent operators to test DL frameworks. The above methods successfully trigger bugs in a single operator, but ignore some bugs which can only be triggered by invoking operator combinations in DL models. To better detect DL framework bugs, some researchers propose methods which regard DL models as test inputs. Among them, Cradle [39] collects publicly available models and directly uses them to test DL frameworks. However, the number of publicly available models is limited. To overcome this challenge, researchers propose various methods to generate test input models. These methods can be mainly categorized into generation-based and mutation-based methods. Generation-based methods construct DL models from scratch. Among them, Gandalf [15] generates new models using context-free grammar and optimizes model generation using deep-Q network. Audee [40] randomly generate a set of DL models by combining different parameter configurations in DL frameworks. Muffin [12] first designs multiple templates, which determine the connections between operators. Then, Muffin generates new models by determining the types of operators in each position of the templates. DLLEN [41], Fuzzgpt [42], and TitanFuzz [43] generate DL models with the aid of LLM (large language model). Mutation-based methods generate new DL models by mutating seed models. Specifically, LEMON [11] chooses publicly available models as seeds, and designs multiple mutation rules to modify the model structure, such as adding operators, deleting operators, and connecting operators. COMET [13] adopts the MH algorithm, a Markov Chain Monte Carlo (MCMC) algorithm to select mutation rules. In this work, we apply a mutation-based method to generate new DL models.
@@
The role of heuristic guidance differs in different methods. Specifically, Muffin [12] adopts heuristic guidance to control both the selection of templates and the selection of specific operators for each position within the templates. LEMON [11] and COMET [13] adopt heuristic guidance to control the selection of mutation seeds and mutation rules. However, existing methods fail to guide the specific implementation of each mutation rule (e.g., controlling the probability of inserting each type of operator). To overcome this challenge, we design multiple-level heuristic guidance for test input model generation, which can not only heuristically control the probability of each mutation rule being selected but also control the specific implementation of the mutation rules.
8 Conclusion
@@
9 Data Availability
Our source code and experimental data are available on: https://github.com/DeeplearningMMM/DLMMM.

### v2

@@ 5.3. RQ4: Effectiveness of Bug Detection
5.3.1. Crashes
As introduced in Section 4.5, DLMMM records the logs while executing test input models to detect crashes. Since one crash may be repeatedly triggered on different datasets, two researchers analyze the logs respectively to remove redundant crashes. Table 4 shows the crashes detected by DLMMM under each dataset. In this table, the first column represents the adopted dataset, and the following three columns represent the number of crashes detected under TensorFlow, PyTorch, and MindSpore. Figure 2 shows the total number of crashes triggered on all datasets after removing redundancy. In this figure, the first column represents the method name, and the following three columns represent the number of crashes detected under each tested DL framework. The experimental results show that DLMMM newly detects 19 crashes, including 3 in TensorFlow, 3 in PyTorch and 13 in MindSpore, which is the most among the four baselines. Specifically, only one crash is triggered by LEMON; 13 are by COMET, and 11 are by Gandalf. Muffin does not trigger any crashes. We have reported all the detected crashes to the public community. Among them, 16 of 19 have been confirmed by developers and two of 19 are fixed in the updated version.
In addition, we analyze the root causes behind detected crashes. In TensorFlow, one crash arises from failure in initialization. Another two crashes are due to unsupported implementation (e.g., unsupported operator sparse_mat_mulsparse\_mat\_mul and unsupported parameter dilationsdilations for the operator $Convolution$). In PyTorch, one crash arises from the failure in creating a descriptor. Another two crashes are due to unsupported implementation. In MindSpore, one crash arises from unsupported implementation in the operator $PRELU$ in the CPU environment. One crash arises from the failure in creating a descriptor. Another 11 crashes are caused by the failure in memory allocation. This issue leads to crashes during executing multiple operators, such as $Convolution$ and Depthwise_ConvolutionDepthwise\_Convolution.
Table 4. Number of Unique Crashes Detected by DLMMM
@@ 5.3.1. Crashes
Figure 2. Number of Unique Crashes Detected by Each Method
5.3.2. NaN & Inconsistency Bugs
Table 5 shows the number of NaN & Inconsistency bugs triggered on all datasets. In this table, the first column represents the adopted dataset, and the following three columns represent the number of NaN & Inconsistency bugs detected under each tested DL framework. This table shows NaN & inconsistency bug is successfully triggered in almost all datasets (except MNIST). The small test input tensor size in MNIST (only 48 in height and width dimension) limits the bug detection effectiveness. Figure 3 shows the number of unique NaN & inconsistency bugs detected by DLMMM and all baselines. In this figure, the first column represents the method name, the following three columns represent the same meaning as Table 5. The experimental results show that DLMMM triggers most NaN & inconsistency bugs compared to four baselines. Specifically, DLMMM triggers 175 NaN & inconsistency bugs in total, including 73 under TensorFlow, 60 under PyTorch, and 42 under MindSpore. LEMON and Muffin do not detect any NaN & inconsistency bugs at all. COMET and Gandalf detect six and 36 NaN & inconsistency bugs, respectively, under all these frameworks. In conclusion, DLMMM performs best in detecting NaN & inconsistency bugs compared to all baselines under all frameworks.
@@ 5.3.2. NaN & Inconsistency Bugs
Table 5. Number of Unique NaN & Inconsistency Bugs Detected by DLMMM
@@
Figure 3. Number of Unique Detected NaN & Inconsistency Bugs
@@
5.4. RQ5: Effectiveness of Test Input Model Generation
@@
5.4.1. Operator Combination Variety
Figure 4 shows the average operator combination variety of DLMMM on each dataset. In this figure, the text on the axis represents the datasets, and the numbers on the bars represent the average operator combination variety of each dataset (measured by variety_degreevariety\_degree introduced in Section 3.2.2). Figure 5 shows the comparison between DLMMM and baselines in average operator combination variety in all datasets. In this figure, the text on the axis represents the method name, and the numbers on the bars represent the average operator combination variety of each method. The experimental results show that the average operator combination variety of test input models generated by DLMMM is 64.25, which greatly exceeds 7.98 in LEMON, 5.51 in Muffin, 18.35 in COMET, and 2.56 in Gandalf. Compared to the best performance of all baselines, DLMMM achieves a 250.13% improvement in operator combination variety.
Figure 4. Operator Combination Variety of DLMMM
Figure 5. Comparison between DLMMM and Baselines in Operator Combination Variety
5.4.2. Model Execution Time (Efficiency)
To evaluate the efficiency of DLMMM, we count the number of models executed by DLMMM in six hours and calculate the average model execution time. The experimental results are shown in Table 6. In this table, the first column represents the adopted dataset. The second column represents the average model execution time. The last column represents the number of executed test input models. In addition, we calculate the average number of executed test input models on all datasets and compare it with four baselines, as shown in Figure 6. In this figure, the first column represents the method name, and the meaning of the following two columns is the same as in Table 6. The experimental results show that, under all datasets, DLMMM takes 5.06 seconds at most and 3.50 seconds on average to execute each single model. However, LEMON, Muffin, COMET, and Gandalf take 63.90 seconds, 14.46 seconds, 122.72 seconds, and 36.92 seconds, respectively. In conclusion, the time consumption of DLMMM to execute a single model is within 35% of all state-of-the-art methods.
Table 6. Efficiency of DLMMM
@@ 5.4.2. Model Execution Time (Efficiency)
Figure 6. Comparison between DLMMM and Baselines in Efficiency
@@
5.5. RQ6: Ablation Study on Heuristic Guidance
@@
We respectively run DLMMM and all baselines for six hours and record detected crashes and NaN & inconsistency bugs. For fairness, all test input tensors are randomly generated and kept the same among DLMMM and all baselines. The experimental results are shown in Table 7. In this table, the first column represents the method name. The following two columns represent the number of detected crashes and NaN & inconsistency bugs. From the experimental results, we have the following conclusions. Firstly, DLMMM (detecting 19 crashes and 175 NaN & inconsistency bugs) detects more bugs than $DLMMM_{r}$ (detecting no crash and one NaN & inconsistency bug), which shows the positive contribution of our heuristic guidance. Secondly, DLMMM detects more bugs than $DLMMM_{p}$ (detecting one crash and eight NaN & inconsistency bugs), $DLMMM_{o}$ (detecting five crashes and two NaN & inconsistency bugs), and $DLMMM_{t}$ (detecting one crash and three NaN & inconsistency bugs), which shows the positive contribution of our measurement fusion. Thirdly, $DLMMM_{p}$, $DLMMM_{o}$ and $DLMMM_{t}$ all detect more bugs than $DLMMM_{r}$, which respectively shows the positive contribution of bug detection effectiveness, operator combination variety, and model execution time.
Table 7. Comparison in Effectiveness in Ablation Study
@@
To investigate whether DLMMM’s designs in heuristic guidance contribute to its testing effectiveness beyond simply generating a larger number of test input models. We record the number of bugs detected by DLMMM when it generates the same number of test input models as existing methods. The experimental results are shown in Table 8. The first column of the table indicates the number of generated test input models. The first four rows show the results when DLMMM generats the same number of models as COMET (176), LEMON (338), Gandalf (585), and Muffin (1494), respectively. The fifth row shows the results using the total number of models generated by DLMMM. The following two columns indicate the number of crashes and NaN & inconsistency bugs detected by DLMMM. As shown in the table, when generating 176 models, DLMMM detects 15 crashes and 22 NaN & inconsistency bugs. In contrast, as shown in Figure 2 and 3, COMET detects only 13 crashes and 6 NaN & inconsistency bugs when generating the same number of models (176). When generating 338 models, DLMMM detects 17 crashes and 30 NaN & inconsistency bugs. However, LEMON detects only 1 crash and 0 NaN & inconsistency bugs with the same model count (338), as shown in Figure 2 and 3. When generating 585 models, DLMMM detects 18 crashes and 43 NaN & inconsistency bugs, whereas Gandalf detects only 11 crashes and 36 NaN & inconsistency bugs under the same condition. When generating 1494 models, DLMMM detects 18 crashes and 66 NaN & inconsistency bugs. However, Muffin does not detect any bugs when generating the same number of models (1494), as shown in Figure 2 and 3. These experimental results demonstrate that DLMMM detects more crashes and NaN & inconsistency bugs than the existing methods when generating the same number of test input models. In summary, besides generating more test input models, the heuristic guidance in DLMMM also contributes to improving testing effectiveness under an equal number of test input models.
Table 8. DLMMM’s Testing Effectiveness When Generating Different Number of Test Input Models
Generated Model Number Crash NaN & Inconsistency
176 15 22
338 17 30
585 18 43
1494 18 66
6177 19 175
Answer to RQ6: The ablation study shows that DLMMM’s heuristic guidance, operator fusion, and each model measurement all contribute to improving DLMMM’s bug detection effectiveness. Besides generating more test input models, DLMMM’s above designs in heuristic guidance also contributes to improving testing effectiveness under an equal number of the test input models.
6. Validity & Threat
@@
The external threat lies in the DL frameworks and operators tested. To eliminate this threat of DL framework selection, we apply our method to test three widely used frameworks, respectively TensorFlow, PyTorch and MindSpore. The other frameworks can also be tested by applying the core idea of our method. Additionally, with the development of large language models (LLMs), developers have encapsulated some commonly used model substructures as operators (e.g., the substructure called Transformer) to facilitate more convenient model construction. To adapt to this trend and enhance the scalability of DLMMM, the measurement of mutation and operator variety in DLMMM is also designed to be applicable to commonly used model substructures (e.g., LSTM and Transformer are also tested operators in DLMMM), rather than simply limited to individual functional operators (such as the operator $Convolution$).
@@ 7. Related Works
DL framework bugs have received extensive attention in recent years. Many empirical studies on DL framework bugs are conducted, which establish the foundation for DL framework testing (DLbug1, ; DLbug2, ; DLbug3, ; DLbug4, ; DLbug5, ; DLbug6, ; DLbug7, ; DLbug8, ; DLbug9, ). DL framework testing basically contains interface-based DL framework testing and model-based DL framework testing. Interface-based DL framework testing separately invokes each operator provided by DL frameworks. Among them, Predoo (predoo, ) mutates test input tensors to maximize output precision error. FreeFuzz (freefuzz, ) conducts automated API-level fuzzing via mining from actual code and model executions in the wild. DocTer (DocTer, ) analyzes API documentation to extract input constraints for each API function provided by DL frameworks. DeepREL (DeepREL, ) infers potential API relations based on their syntactic or semantic information, and synthesizes test programs for invoking relational APIs. Duo (duo, ) generates test input tensors by employing nine mutation rules derived from genetic algorithms. EAGLE (eagle, ) invokes equivalent operators to test DL frameworks. The above methods successfully trigger bugs in a single operator, but ignore some bugs which can only be triggered by invoking operator combinations in DL models. To better detect DL framework bugs, some researchers propose methods which regard DL models as test inputs. Among them, Cradle (cradle, ) collects publicly available models and directly uses them to test DL frameworks. However, the number of publicly available models is limited. To overcome this challenge, researchers propose various methods to generate test input models. These methods are mainly categorized into generation-based and mutation-based methods. Generation-based methods construct DL models from scratch. Among them, Gandalf (gandalf, ) generates new models using context-free grammar and optimizes model generation using deep-Q network. Audee (audee, ) randomly generates a set of DL models by combining different parameter configurations in DL frameworks. Muffin (muffin, ) first designs multiple templates, which determines the connections between operators. Then, Muffin generates new models by determining the types of operators in each position of the templates. DLLEN (dllen, ), Fuzzgpt (fuzzgpt, ), and TitanFuzz (titanfuzz, ) generate DL models with the aid of large language models. Mutation-based methods generate new DL models by mutating seed models. Specifically, LEMON (lemon, ) chooses publicly available models as seeds, and designs multiple mutation rules to modify the model structure, such as adding operators, deleting operators, and connecting operators. ModelMeta (ModelMeta, ) designs Metamorphic Relations (MRs) based on input data and parameter settings to generate equivalent test input models. DevMuT (mu2024devmut, ) designs model mutation rules based on the developer expertise. DLJSFuzzer (DLJSFuzzer, ) designs mutation rules to generate JavaScript test input models. COMET (comet, ) adopts the MH algorithm, a Markov Chain Monte Carlo (MCMC) algorithm to select mutation rules. In this work, we apply a mutation-based method to generate DL models.
@@
The role of heuristic guidance differs in different methods. Specifically, Muffin (muffin, ) adopts heuristic guidance to control both the selection of templates and the selection of specific operators for each position within templates. LEMON (lemon, ) and COMET (comet, ) adopt heuristic guidance to control the selection of mutation seeds and mutation rules. However, existing methods fail to guide the specific implementation of each mutation rule (e.g., controlling the probability of inserting each type of operator). To overcome this challenge, we design multiple-level heuristic guidance for test input model generation, which can not only heuristically control the probability of each mutation rule being selected but also control the specific implementation of the mutation rules.
8. Conclusion